OpenStudy (anonymous):

how can I prove that d/dx (1/y) = 0. ??? how can I prove that d/dx (1/y) = 0. ??? @MIT 18.01 Single …

OpenStudy (anonymous):

d/dx of 1/y = -1/y2 not 0, I do not know if you can prove it to be 0.

OpenStudy (anonymous):

why is it occurring?

OpenStudy (anonymous):

if I put d/dx of 1/y in MS math or wolfram it return 0

OpenStudy (jamesj):

First of all \[ \frac{d \ }{dx} (1/x) = -\frac{1}{x^2} \] Now, if y is a function of x, y = y(x), then using implicit differentiation and the chain rule, \[ \frac{d \ }{dx} (1/y) = -\frac{1}{y^2} \frac{dy}{dx} \] But if \(y\) is just a constant, completely unrelated to \(x\), then \(1/y\) is also just a constant (assuming of course that \(y\) is not zero), and for any constant \(c\) \[ \frac{d \ }{dx} c = 0 \]

OpenStudy (jamesj):

The reason Wolfram is giving you that (d/dx) of 1/y is zero is it is treating y as a constant, unrelated to x. That may be correct, depending on your context. But very often we are treating y as the vertical graphing variable and y is dependent on x, as y = f(x). In this case, the derivative is most emphatically not zero, as I just showed above.